# A Study of Undefined Behavior Across Foreign Function Boundaries in Rust Libraries

### v5

@@ IV-B Typing
We found 11 bugs in 10 crates involving typing or initialization. Most were caused by incorrect use of MaybeUninit<T>, but several involved incorrect foreign function bindings.
@@
We found four bugs that involved partial initialization of MaybeUninit<T>. Bugs #0 and #0 were found in the same library. In each case, an uninitialized instance of a struct T containing an array of values was created using MaybeUninit<T> and passed across the FFI to be initialized. The foreign call only initialized the first element of the array to zero. The remaining uninitialized elements would never be read by C, since every iteration stopped at the null terminator. Miri reported these initialization patterns as an error when MaybeUninit<T>::assume_init() was called. We fixed each bug by zero-initializing the entire array in Rust.
Bug #0 seemed to involve an equivalent pattern; MaybeUninit<T>::assume_init() triggered an error for a struct that was initialized by a foreign function. However, the function properly zero-initialized the struct using memset. After examining the LLVM bytecode of the foreign function, we found that all 88 bytes of the struct were initialized on the LLVM stack, but only the first 80 bytes were copied back into Rust. These missing bytes corresponded to padding fields, which were never directly accessed by the foreign library. Optimizations had been disabled, but LLVM still removed this unnecessary write access. We fixed Bug #0 by zero-initializing the padding fields before calling assume_init().
@@
We found six crates with one or more incorrect foreign function bindings. All of these incorrect bindings had been written manually. Three of these crates had bindings with missing return types. In Bug #0, a binding was declared in a unit test without its 32-bit integer return type. The C implementation used this to return a status code indicating if one of its integer parameters was within bounds. The remaining two bugs involved bindings with incorrect integer types, which led to incomplete initialization. In Bug #0, a function was declared to return a 32-bit integer, but its C implementation returned a boolean value. In Bug #0, the last parameter of a function was declared as a 32-bit integer, but the C implementation expected size_t. The width of size_t is architecture-dependent, so this binding was only correct for 32-bit architectures. To fix this bug, multiple prior releases needed to be patched to use Rust’s usize type, which has similar semantics.
@@ IV-C Allocation
We found 12 issues related to allocation in 12 distinct crates. These include 10 memory leaks and one new cross-language deallocation bug. We also found one null pointer dereference, but it had no connection to Rust.
@@
Rust’s Box<T> and CString use similar APIs to encapsulate heap allocations. For each type, the function into_raw consumes an instance and produces a pointer to its interior heap allocation. To avoid a memory leak, this pointer must be reconstituted in its wrapper type using the function from_raw. We found four leaks caused by calling into_raw on either Box<T> or CString without later calling from_raw. Each type was used to allocate memory for a foreign function call, and each bug was fixed by adding a call to from_raw after the function returned. The remaining six leaks were of memory that had been allocated by C. Most were caused by neglecting to call a destructor that had been exposed by the C API. However, in Bug #0, the C API did not expose a destructor, since it had been designed based on the assumption that users would be able to call free. We could not use Rust’s std::alloc::dealloc to fix this, since that could lead to invalid cross-language deallocation.
@@
It is potentially undefined behavior to free a pointer allocated by C in Rust, or vice versa, since each language’s binary may use a different allocator. We found one example of invalid cross-language deallocation. In Bug #0, a pointer to a heap-allocated string was returned to Rust by a foreign function call and stored as a Cow<&’static [u8]>, which will lazily clone its data when mutated. This wrapper type is an enumeration with two variants; Cow::Borrowed receives a reference to a value, while Cow::Owned takes ownership of a value, deallocating it when it goes out of scope. When this Cow was dropped, it deallocated the C heap memory using Rust’s allocator. There was not an immediate fix for this bug, since the C API did not expose a destructor.
@@ VII Threats to Validity
Crates that could not compile with our nightly toolchain were excluded from our evaluation, even if they would have otherwise been compatible with MiriLLI. Limited support for certain features, such as dynamically linked libraries, also prevented us from achieving high test coverage; 61% of tests terminated due to an unsupported operation and 10% timed out. Additionally, the majority of libraries we examined linked against C, and not C++. Crates that failed to compile using our nightly toolchain were excluded from the evaluation, as well as crates that overrode the system’s default C and C++ compiler. However, these limitations did not prevent us from answering our research questions. Our goal was not to create a production-ready tool, but to understand the types of errors that can occur in Rust libraries that call foreign functions. The bugs that occur in general may differ in unexpected ways from those found with our experimental setup.
@@ VIII Conclusion
We conducted a large-scale evaluation to investigate the types of undefined behavior that occur in Rust libraries that call foreign functions. We created MiriLLI---a dynamic analysis tool that combines two existing interpreters---and we used it to identify 47 new bugs from 37 libraries. Though many of these bugs were aliasing violations, Rust’s Tree Borrows aliasing model was more permissive in this context than the earlier Stacked Borrows model. To ensure that these errors are easy to detect, the Rust project should invest in new dynamic analysis tools that can accommodate multi-language applications.

### v6

@@ IV-B Typing
We found 12 bugs from 12 crates that involved typing or initialization. Some were caused by incorrect use of MaybeUninit<T>, while others involved incorrect foreign function bindings.
@@
We found five bugs that involved partial initialization. In Bug #0, an uninitialized instance of a struct T containing an array of values was created using MaybeUninit<T> and passed across the FFI to be initialized. The foreign call only initialized the first element of the array to zero. The remaining uninitialized elements would never be read by C, since every iteration stopped at the null terminator. Miri reported these initialization patterns as an error when MaybeUninit<T>::assume_init was called. We fixed this bug by zero-initializing the entire array in Rust.
Bug #0 seemed to involve an equivalent pattern; MaybeUninit<T>::assume_init triggered an error for a struct that was initialized by a foreign function. However, the function properly zero-initialized the struct using memset. After examining the LLVM bytecode of the foreign function, we found that all 88 bytes of the struct were initialized on the LLVM stack, but only the first 80 bytes were copied back into Rust. These missing bytes corresponded to padding fields, which were never directly accessed by the foreign library. Optimizations had been disabled, but LLVM still removed this unnecessary write access. We fixed Bug #0 by zero-initializing the padding fields before calling assume_init.
@@
We found six crates with one or more incorrect foreign function bindings. All of these incorrect bindings had been written manually. Three of these crates had bindings with missing return types. In Bug #0, a binding was declared in a unit test without its 32-bit integer return type. The C implementation used this to return a status code indicating if one of its integer parameters was within bounds. The remaining three bugs involved bindings with incorrect integer types, which led to incomplete initialization. In Bug #0, a function was declared to return a 32-bit integer, but its C implementation returned a boolean value. In Bug #0, the last parameter of a function was declared as a 32-bit integer, but the C implementation expected size_t. The width of size_t is architecture-dependent, so this binding was only correct for 32-bit architectures. To fix this bug, multiple prior releases needed to be patched to use Rust’s usize type, which has similar semantics.
@@ IV-C Allocation
We found 10 issues related to allocation in 10 distinct libraries. These include nine memory leaks and one new cross-language deallocation bug.
@@
Rust’s Box<T> and CString use similar APIs to encapsulate heap allocations. For each type, the function into_raw consumes an instance and produces a pointer to its interior heap allocation. To avoid a memory leak, this pointer must be reconstituted in its wrapper type using the function from_raw. We found four leaks caused by calling into_raw on either Box<T> or CString without later calling from_raw. Each type was used to allocate memory for a foreign function call, and each bug was fixed by adding a call to from_raw after the function returned. The remaining five leaks were of memory that had been allocated by C. Most were caused by neglecting to call a destructor that had been exposed by the C API. However, in Bug #0, the C API did not expose a destructor, since it had been designed based on the assumption that users would be able to call free. We could not use Rust’s allocator API to fix this, since that could lead to invalid cross-language deallocation.
@@
It is potentially undefined behavior to free a pointer allocated by C in Rust, or vice versa, since each language’s binary may use a different allocator. We found one new example of cross-language deallocation. In Bug #0, a pointer to a heap-allocated string was returned to Rust by a foreign function call and stored as a Cow<&’static [u8]>, which will lazily clone its data when mutated. This wrapper type is an enumeration with two variants; Cow::Borrowed receives a reference to a value, while Cow::Owned takes ownership of a value, deallocating it when it goes out of scope. When this Cow was dropped, it deallocated the C heap memory using Rust’s allocator. There was no immediate fix for this bug, since the C API did not expose a destructor.
@@ VII Threats to Validity
Several of our bug reports are still pending responses, so we cannot be entirely certain that all of our results are novel or true positives. However, we did exclude bugs from our evaluation that had already been reported. We did not independently confirm that these bugs were absent from test cases that did not call foreign functions, but we have no reason to believe that any would have been detected by an unmodified version of Miri at the time of our evaluation.
@@
Crates that could not compile with our nightly toolchain were excluded from our evaluation, even if they would have otherwise been compatible with MiriLLI. Limited support for certain features, such as dynamically linked libraries, also prevented us from achieving high test coverage; 61% of tests terminated due to an unsupported operation and 10% timed out. Additionally, the majority of libraries we examined linked against C, and not C++. Crates that failed to compile using our nightly toolchain were excluded from the evaluation, as well as crates that overrode the system’s default C and C++ compiler. However, these limitations did not prevent us from answering our research questions. Our goal was not to create a production-ready tool, but to understand the types of errors that can occur in Rust libraries that call foreign functions. The bugs that occur in general may differ in unexpected ways from those found with our setup.
@@ VIII Conclusion
We conducted a large-scale evaluation to investigate the types of undefined behavior that occur in Rust libraries that call foreign functions. We created MiriLLI---a dynamic analysis tool that combines two existing interpreters---and we used it to find 46 bugs in 37 libraries. Though many of these bugs were aliasing violations, Rust’s Tree Borrows aliasing model was more permissive in this context than the earlier Stacked Borrows model. To ensure that these errors are easy to detect, the Rust project should invest in new dynamic analysis tools that can accommodate multi-language applications.